# Photonic Doping of Epsilon-Near-Zero Bragg Microcavities

### v1

@@ 3. Numerical Study of Photonic Doping in all-dielectric ENZ Cavities
To numerically investigate these effects and the photonic doping features of ENZ cavities, we examine the resonance behavior of an array of infinitely long dielectric cylinders embedded in ENZ BR cavities. Two-dimensional Comsol simulations are performed to calculate the transmittance of the structure shown in Fig. 4, and the eigenmodes are obtained using the mode analysis module.
@@
Figure 4: Simulation domain of the ENZ Bragg cavity loaded with an array of infinitely-long dielectric cylinders of radius $R=150~\mathrm{nm}$ and refractive index $n=3.5$.
1.1 3.1. Excitation of Transverse Electric (TE) Resonances
@@
Figure 5 presents the transmittance spectra of a cavity, both with and without the NC array in the core. The cavity consists of a half-wave SiO2 core with refractive index of $n_{\mathrm{SiO_{2}}}=1.44+i10^{-6}$ 32, sandwiched between 10 pairs of quarter-wave cladding layers of SiO2 and SiN ($n_{\mathrm{SiN}}=2+i10^{-7}$)33. The cutoff wavelength of the cavity is set to $\lambda_{c}=1500~\mathrm{nm}$. We dope the system with an array of silicon NCs with radii $R=150~\mathrm{nm}$ and refractive indices $n_{\mathrm{Si}}=3.5+i1.38\times 10^{-13}$ 34, positioned at the center of the core. The periodicity of the array is $p=500~\mathrm{nm}$ and $p=700~\mathrm{nm}$ in Figs. 5b and 5c, respectively.
@@
Figure 5: Transmittance spectra and corresponding field profiles: (a) bare Bragg cavity (cutoff wavelength $\lambda_{c}=1500\,\mathrm{nm}$) consisting a half-wave SiO2 core, sandwiched between 10 pairs of quarter-wave cladding layers of SiO2 and SiN; (b,c) the same cavity loaded with arrays of infinitely-long silicon NCs, placed at the core center, with lattice periods (b) $p=500\,\mathrm{nm}$ and (c) $p=700\,\mathrm{nm}$. All structures are illuminated by a normally incident plane wave (along $y$-axis) with the magnetic field $H$ along $z$-axis (TE, $H_{z}$).
@@
Figure 5a shows the transmittance spectrum and field profiles of the bare cavity mode, with a transmission peak at $\lambda=1500~\mathrm{nm}$. The electric field is mainly concentrated at the core center, while the magnetic field is localized at the core boundaries. Introducing the NC array (Fig. 5b) gives rise to two resonant modes within the Bragg stop band, one at a longer wavelength and the other at a shorter wavelength relative to the bare-cavity cutoff. As demonstrated later, these correspond to NZI modes of the structure. The higher-energy mode, located at $\lambda=1371~\mathrm{nm}$, is dominantly confined inside the cylinders and can be identified as the Mie TE0-like mode, associated with the magnetic dipole resonance of the Mie coefficient $a_{0}$. This mode exhibits an enhanced quality factor of $Q=626$, nearly $50$ times higher than its free-space resonance. The lower-energy NZI mode, at $\lambda=1623.5~\mathrm{nm}$ with a quality factor of $Q=607$, is characterized by distinct electric and magnetic field distributions that are mainly localized outside the cylinders. Increasing the periodicity of the array preserves the general field distribution patterns of both the high- and low-energy modes. However, the $Q$-factors of the modes increase, and as shown in Fig. 5c, the lower-energy mode exhibits a blue shift while the higher-energy mode undergoes a red shift, which can extend beyond the cutoff wavelength of the bare cavity (see Fig. 6).
@@
By increasing the periodicity to $p=900~\mathrm{nm}$, the quality factor of the TE0-like mode rises to $Q=2.32\times 10^{4}$, which is about $2\times 10^{3}$ times higher than its free-space Mie counterpart and more than 14 times higher than that of the bare cavity. As shown in Fig.6, additional resonant structural modes also appear, with electromagnetic fields predominantly distributed within the Bragg layers but exhibiting exceptionally high quality factors; for example, $Q=4.87\times 10^{5}$ at $\lambda=1459.1~\mathrm{nm}$, more than 21 times higher than the confined TE0-like mode.
Figure 6: Transmittance spectrum and field profiles of the resonant modes in the same Bragg cavity of Fig. 5b, but with the period of $p=900~\mathrm{nm}$, illuminated by an incident TE plane wave with the magnetic field $H$ normal to the image plane.
@@
Comparison of the magnetic field distribution of the TE0-like mode with that of the bare-cavity mode (Fig. 5a) reveals that enhanced field localization, together with the significant increase in $Q$-factor, produces a magnetic hot spot within the NC, which is one of the key features of photonic doping in ENZ media 10.
@@
Figure 7: Transmittance spectra of a single NC of radius $R=150~\mathrm{nm}$ (solid curve) in a SiO2 matrix as well as arrays of such NCs with periods varying between 400-900 nm.
In these structures, the periodicity of the NC array plays a crucial role in achieving high-$Q$ resonances. However, the effect of periodicity on the transmission peak of the same array in free space is opposite. To demonstrate this, we calculated the transmittance of a plane wave for a single NC and for arrays of identical NCs with varying periodicity. The results show that arranging the NCs into an array causes the transmission peak to broaden and shift to shorter wavelengths, with the effect becoming more pronounced at smaller periods.
1.2 3.1.1 PMC and EMNZ Features in Photonically Doped BR Cavities
@@
To demonstrate the PMC and EMNZ features by photonic doping of BR cavities, the same structure as in Fig. 5b is considered, but with increased number of Bragg layers to $N=14$ pairs of quarter-wave layers. The transmittance spectrum of the structure is shown in Fig. 8a with the corresponding field profiles of the TE0-like mode at $\lambda=1372.38~\mathrm{nm}$. We have also calculated the dispersive effective refractive index (using the mode analysis module of COMSOL), as well as the effective relative permittivity and permeability around this mode. The results are shown in Fig. 8b.
@@
Figure 8: (a) Transmittance and resonant-mode field distribution corresponding to a BR cavity with $N=14$ pairs of cladding layers on each side of the half-wave core and doped with an array of cylindrical NCs of radius $R=150~\mathrm{nm}$ having a period of $p=500~\mathrm{nm}$. (b) Calculated effective refractive index as well as effective relative permittivity and permeability around the wavelength $\lambda=1372.38~\mathrm{nm}$ corresponding to the TE0-like resonant mode in the same NC-doped BR cavity.
@@
In order to consistently define the effective magnetic permeability of a periodic array of inclusions, we follow the homogenization procedure introduced by Silveirinha and Engheta 35. For a periodic Bloch mode, the microscopic electric and polarization current density fields have the form 36
@@
where $\tilde{\mathbf{E}}$ and $\tilde{\mathbf{J}}$ are periodic functions over the unit cell and $\mathbf{k}_{B}$ is the Bloch wavevector. Accordingly, the microscopic magnetic dipole moment per unit length is written as 36
$m_{z}\;=\;\frac{1}{2}\iint_{\text{inc}}\Big[(x-x_{c})J_{y}-(y-y_{c})J_{x}\Big]\,e^{-i\mathbf{k}_{B}\cdot\mathbf{r}}\,dA,$ (9)
@@
$M_{z}=\frac{m_{z}}{A_{\text{cell}}},\qquad A_{\text{cell}}=\iint_{\text{cell}}dA.$ (10)
@@
Since the direct cell average of the magnetic field, $\langle H_{z}\rangle$, contains contributions from both the applied field and the self-field of the induced dipole, a corrected bulk magnetic field is introduced as 5
@@
The TE0-like mode resonates at $\lambda=1372.38~\mathrm{nm}$, in the close vicinity of the cutoff. The PMC-like response appears at the wavelength where the effective permeability (solid green) exhibits a pole, $\mu_{r}\to\infty$, corresponding to a high-impedance condition. The EMNZ region, corresponding to wavelengths shorter than the cutoff wavelength, is highlighted in blue. In this region, both the effective permittivity and permeability are positive and smaller than unity. For wavelengths longer than the cutoff wavelength, the signs of the effective permittivity and permeability are opposite, and consequently the wave cannot propagate in this spectral range, as indicated by the rapid growth of the imaginary part of the effective index (red-circle curve).
@@
To excite the TM resonances of the doped BR cavity, it is illuminated by a normally incident plane wave with its electric field polarized along the $z$-axis. The transmittance and field profile of the TM1-like mode are shown in Fig. 9. For a period of $p=500~\mathrm{nm}$, Fig. 9a demonstrates that only the TM1-like mode exists, which is confined inside the NCs. In contrast, Fig. 9b, corresponding to $p=750~\mathrm{nm}$, reveals the emergence of an additional mode at a shorter wavelength, primarily confined in the regions between the NCs. When the period is further increased to $p=900~\mathrm{nm}$, both modes undergo a red shift, exhibit an enhanced $Q$-factor, and become spectrally closer to each other. At the same time, other modes also appear at shorter wavelengths, with their fields predominantly distributed in the Bragg mirror layers rather than within the cavity core.
@@
Figure 9: Transmittance and the resonant-mode field distribution of a BR cavity with the same parameters as in Fig. 5, loaded with an array of cylindrical NPs of radius $R=150~\mathrm{nm}$, for periods of (a) $p=500~\mathrm{nm}$; (b) $p=750~\mathrm{nm}$; and (c) $p=900~\mathrm{nm}$. All structures are illuminated by a normally incident plane wave (along $y$-axis) with the electric field $E$ along $z$-axis (TM, $E_{z}$).
1.4 3.3 Tunability of the Hybrid Bragg-Mie Modes Beyond the Size Parameter
@@
In addition to the conventional size parameters, $x_{i}=2\pi n_{i}r/\lambda$ and $x_{e}=2\pi n_{e}r/\lambda$, which depend on wavelength, particle size and the refractive indices of the particle and surrounding medium,13 the hybrid Bragg–Mie modes can be further tuned by external and structural parameters. Specifically, varying the angle of incidence and the bare-cavity cutoff wavelength introduces additional means for controlling the resonance wavelength of the Bragg-Mie modes. These parameters enable tunability of the hybrid resonances beyond what is achievable with the size parameter alone, offering a richer degree of freedom for engineering their spectral position and field confinement. For instance, this approach can enable straightforward fine tuning and alignment of the mode resonances with the magnetic dipole resonances of doped atoms or quantum dots, facilitating enhancement or selective excitation of their intrinsic magnetic dipole transitions.
@@
Figure 10 illustrates this tunability for both the TE0-like (Fig. 10a) and TM1-like (Fig. 10b) modes by showing the transmittance of a doped BR cavity with the same parameters as in Fig. 5, for $p=500~\mathrm{nm}$ and different bare-cavity cutoff wavelengths of $\lambda_{c}=1450~\mathrm{nm}$, $\lambda_{c}=1500~\mathrm{nm}$, and $\lambda_{c}=1550~\mathrm{nm}$. In both cases, the resonance exhibits a blueshift for smaller cutoff wavelengths and a redshift for larger cutoff wavelengths.
@@
Figure 10: Transmittance spectra of a doped BR cavity with $p=500~\mathrm{nm}$ and different cutoff wavelengths $\lambda_{c}=1450~\mathrm{nm}$ (red dashed), $\lambda_{c}=1500~\mathrm{nm}$ (blue solid), and $\lambda_{c}=1550~\mathrm{nm}$ (green dotted). (a) TE0-like mode, (b) TM1-like mode.
@@
Figure 11 shows the TE-polarized transmittance of the structure in Fig. 10 for increasing AoI. The lower- and higher-energy modes shift and move closer together, illustrating an angle-tunable response; the lower-energy branch blue-shifts across the range, while the higher-energy branch shows a weaker, non-monotonic shift at larger angles.
Figure 11: TE-polarized transmittance of a structure with the same parameters of Fig. 10, and a cutoff wavelength of $1500~\mathrm{nm}$. Results are shown for incidence angles of $0^{\circ}$ (solid curve), $25^{\circ}$ (dashed curve), and $50^{\circ}$ (dash-dotted curve).
@@
Figure 12 shows the TM-polarized transmittance of the same structure as in Fig. 11, but with $p=750~\mathrm{nm}$ and AoIs of $0^{\circ}$, $5^{\circ}$, and $15^{\circ}$. As seen in Fig. 12a, for nonzero AoIs a new resonance emerges with a markedly higher $Q$-factor, particularly at smaller angles (e.g., $Q\approx 1.5\times 10^{4}$ for AoI $5^{\circ}$, compared to $Q\approx 700$ for the TM1-like mode). This resonance is structurally similar to the TM1-like mode but exhibits a $90^{\circ}$ rotated field distribution. With increasing AoI, the new mode undergoes a red shift, while the TM1-like mode gradually vanishes. As illustrated in Fig. 12b for an AoI of $15^{\circ}$, additional resonances also appear at shorter wavelengths, with their fields predominantly distributed in the Bragg mirror layers rather than in the cavity core.
@@
Figure 12: TM-polarized transmittance of a BR cavity with $N=10$ pairs of cladding layers on each side of a half-wave core, designed for a cutoff wavelength of $1500~\mathrm{nm}$, and loaded with an array of cylindrical NPs of radius $R=150~\mathrm{nm}$ and period $p=750~\mathrm{nm}$. (a) Results for incidence angles of $0^{\circ}$ (dashed curve) and $5^{\circ}$ (solid curve). (b) Results for an incidence angle of $15^{\circ}$.
@@
The modes discussed above, including the TE- and TM-like resonances confined inside the NCs, the modes localized between the NCs, and the new high-$Q$ resonance emerging at nonzero AoI in Fig. 12a, can all be identified as high-FoM NZI modes of the doped BR cavity. This is confirmed by the dispersion curves shown in Fig. 13a in the wavelength range of 1300-1650 nm, together with their corresponding transmission peaks presented in Fig. 13b.
@@
Figure 13: (a) Real and imaginary parts of the effective refractive index of TE and TM modes, together with their corresponding field distributions, in a BR cavity with $N=14$ pairs of cladding layers on each side of a half-wave core, loaded with an array of cylindrical NPs of radius $R=150~\mathrm{nm}$ and refractive index $n=3.5$, with a period of $p=750~\mathrm{nm}$. (b) Transmission spectrum of the same cavity for TE (solid curve) and TM (dashed/dotted curves) modes.

### v2

@@ 3 Numerical Study of Photonic Doping in ENZ Bragg Cavities
To numerically investigate these effects and the photonic doping features of ENZ cavities, we examine the resonance behavior of a periodic array of infinitely long dielectric cylinders embedded in ENZ Bragg cavities. Two-dimensional Comsol simulations are performed to calculate the transmittance of the structure shown in Fig. 4, and the eigenmodes are obtained using the mode analysis module.
@@
Figure 4: Simulation domain of the ENZ Bragg cavity loaded with a periodic array of infinitely-long silicon cylinders of radius $R=150~\mathrm{nm}$. The domain is enclosed by periodic and scattering boundary conditions (BCs).
3.1 Excitation of Transverse Electric Resonances
@@
Transverse Electric (TE) resonances arise when the electric field component of the incident wave is oriented perpendicular to the NC axis. Figure 5 presents the transmittance spectra of a cavity, both with and without the NC array in the core. The cavity consists of a half-wave SiO2 core with refractive index of $n_{\mathrm{SiO_{2}}}=1.44+i10^{-6}$ [Malitson1965], sandwiched between two Bragg mirrors, each composed of 10 pairs of quarter-wave cladding layers made of SiO2 and SiN ($n_{\mathrm{SiN}}=2+i10^{-7}$)[Beliaev2022]. The cutoff wavelength of the cavity is set to $\lambda_{c}=1500~\mathrm{nm}$. The system is doped with an array of silicon NCs with radii $R=150~\mathrm{nm}$ and refractive index $n_{\mathrm{Si}}=3.5+i1.38\times 10^{-13}$ [Polyanskiy2024], positioned at the center of the core. The periodicity of the array is $p=500~\mathrm{nm}$ and $p=700~\mathrm{nm}$ in Figs. 5b and 5c, respectively.
@@
Figure 5: Transmittance spectra and corresponding field profiles. (a) Bare Bragg cavity (cutoff wavelength $\lambda_{c}=1500\,\mathrm{nm}$) consisting a half-wave SiO2 core, sandwiched between 10 pairs of quarter-wave cladding layers of SiO2 and SiN. (b,c) The same cavity loaded with arrays of infinitely-long silicon NCs, placed at the core center, with lattice periods (b) $p=500\,\mathrm{nm}$ and (c) $p=700\,\mathrm{nm}$. All structures are illuminated by a normally incident plane wave (along $y$-axis) with the magnetic field $H$ along $z$-axis (TE, $H_{z}$).
@@
Figure 5a shows the transmittance spectrum and field profiles of the bare cavity mode, with a transmission peak very close to the cutoff wavelength, $\lambda_{c}=1500\,\mathrm{nm}$. The electric field is mainly concentrated to the center of the core, while the magnetic field is localized at the core boundaries. Introduction of the NC array (Fig. 5b) turns the bare cavity mode into two resonant modes within the Bragg stop band, one at a longer wavelength and the other at a shorter wavelength relative to the bare-cavity cutoff. As demonstrated later, these correspond to NZI modes of the structure. The higher-energy mode, located at $\lambda=1371~\mathrm{nm}$, is dominantly confined inside the cylinders and can be identified as the Mie TE0-like mode, associated with the magnetic dipole resonance of the Mie coefficient $a_{0}$. This mode exhibits an enhanced quality factor of $Q=626$, nearly $50$ times higher than its free-space resonance. The lower-energy NZI mode, at $\lambda=1623.5~\mathrm{nm}$ with a quality factor of $Q=607$, is characterized by distinct electric and magnetic field distributions that are mainly localized outside the cylinders. Increasing the periodicity of the array preserves the general field distribution patterns of both the high- and low-energy modes. However, the $Q$-factors of the modes increase, and as shown in Fig. 5c, the lower-energy mode exhibits a blue shift while the higher-energy mode undergoes a red shift, which can extend beyond the cutoff wavelength of the bare cavity (see Fig. 6).
@@
By increasing the periodicity to $p=900~\mathrm{nm}$, the $Q$-factor of the TE0-like mode rises to $Q=2.32\times 10^{4}$, which is about $10^{3}$ times higher than for corresponding free-space array and over 14 times higher than that of the bare cavity. As shown in Fig. 6, additional resonant structural modes also appear, with electromagnetic fields predominantly distributed within the Bragg layers but exhibiting exceptionally high quality factors; for example, $Q=4.87\times 10^{5}$ at $\lambda=1459.1~\mathrm{nm}$, even higher than that of the confined TE0-like mode.
Figure 6: Transmittance spectrum and field profiles of the resonant modes in the same Bragg cavity as in Fig. 5b, with a period of $p=900\,\mathrm{nm}$, under TE plane-wave illumination ($\mathbf{H}\parallel z$). To visually resolve closely spaced resonances, two wavelength intervals highlighted in blue are displayed with nonuniform horizontal scaling: the regions $1335\text{--}1345\,\mathrm{nm}$ and $1355\text{--}1365\,\mathrm{nm}$ are expanded by factors of 4 and 8, respectively, relative to the rest of the spectrum.
@@
Comparison of the magnetic field distribution of the TE0-like mode with that of the bare-cavity mode (Fig. 5a) reveals that enhanced field localization, together with the significant increase in $Q$-factor, produces a magnetic hot spot within the NCs, which is one of the key features of photonically doped ENZ media [Liberal2017c].
In these structures, although the period of the doped NC array modulates the transmission spectrum, leading to shifts in the resonance position and changes in bandwidth, the array periodicity by itself, when decoupled from the cavity, does not support the formation of narrowband resonances. This indicates that the emergence of high-Q resonances is primarily governed by the ENZ cavity rather than by collective lattice effects. This conclusion is supported by Fig. 7, which compares the transmittance of a TE-polarized plane wave through a single NC with that through arrays of identical NCs with varying periodicities. The curves show that arranging the NCs into an array causes the transmission peak to broaden and shift to shorter wavelengths, with the effect becoming more pronounced at smaller periods.
@@
Figure 7: Transmittance spectra of a single NC of radius $R=150~\mathrm{nm}$ (solid curve) and index $n_{\mathrm{Si}}$ in a SiO2 matrix as well as arrays of such NCs with periods varying between 400–900 nm.
3.1.1 PMC and EMNZ Features in Photonically Doped Cavities
@@
To demonstrate the PMC and EMNZ features by photonic doping of Bragg cavities, the same structure as in Fig. 5b is considered, but with increased number of Bragg layers to $N=14$ pairs of quarter-wave layers. The transmittance spectrum of the structure is shown in Fig. 8a, together with the corresponding field profiles of the TE0-like mode at $\lambda=1372.38~\mathrm{nm}$. We also computed the dispersive effective refractive index (using COMSOL’s mode analysis module), along with the effective relative permittivity and permeability around this mode, as presented in Fig. 8b.
@@
Figure 8: (a) Transmittance and resonant-mode field distribution corresponding to a Bragg cavity with $N=14$ pairs of cladding layers on each side of the half-wave core and doped with an array of cylindrical NCs of radius $R=150~\mathrm{nm}$ having a period of $p=500~\mathrm{nm}$. (b) Calculated effective refractive index as well as effective relative permittivity and permeability around the wavelength $\lambda=1372.38~\mathrm{nm}$ corresponding to the TE0-like resonant mode in the same NC-doped Bragg cavity.
@@
In order to consistently define the effective magnetic permeability of a periodic array of inclusions, we follow the homogenization procedure introduced by Silveirinha and Engheta [Silveirinha2006]. For a periodic Bloch mode, the microscopic electric and polarization current density fields have the form [Chremmos2015]
@@
where $\tilde{\mathbf{E}}$ and $\tilde{\mathbf{J}}$ are periodic functions over the unit cell and $\mathbf{k}_{B}$ is the Bloch wavevector. Accordingly, the microscopic magnetic dipole moment per unit length is written as [Chremmos2015]
$m_{z}\;=\;\frac{1}{2}\iint_{\text{inc}}\Big[(x-x_{c})J_{y}-(y-y_{c})J_{x}\Big]\,e^{-i\mathbf{k}_{B}\cdot\mathbf{r}}\,\mathrm{d}A,$ (9)
@@
$M_{z}=\frac{m_{z}}{A_{\text{cell}}},\qquad A_{\text{cell}}=\iint_{\text{cell}}\mathrm{d}A.$ (10)
@@
Since the direct cell average of the magnetic field $\langle H_{z}\rangle$ contains contributions from both the applied field and the self-field of the induced dipole, a corrected bulk magnetic field is introduced as [Silveirinha2007]
@@
The TE0-like mode resonates at $\lambda=1372.38~\mathrm{nm}$, in the close vicinity of the cutoff. The PMC-like response appears at the wavelength where the effective permeability (solid green) exhibits a pole ($\mu_{r}\to\infty$), corresponding to a high-impedance condition. The EMNZ region, corresponding to wavelengths shorter than the cutoff wavelength, is highlighted in blue. In this region, both the effective permittivity and permeability are positive and smaller than unity. For wavelengths longer than the cutoff wavelength, the signs of the effective permittivity and permeability are opposite, and consequently the wave cannot propagate in this spectral range, as indicated by the rapid growth of the imaginary part of the effective index (red-circle curve).
@@
The TM resonances of the doped Bragg cavity are excited by illumination with a normally incident plane wave with its electric field polarized along the $z$-axis. The transmittance and field profiles of TM1-like modes are shown in Fig. 9. For a period of $p=500~\mathrm{nm}$ only the TM1-like mode exists, and is confined inside the NCs (Fig. 9a). In contrast, when the period is increased to $p=750~\mathrm{nm}$, an additional mode emerges at shorter wavelengths (Fig. 9b). The fields of this mode are primarily confined to the regions between the NCs. When the period is further increased to $p=900~\mathrm{nm}$, both modes undergo a redshift, approach each other, and exhibit enhanced $Q$-factors (Fig. 9c). At the same time, other modes also appear at shorter wavelengths, with their fields predominantly distributed in the Bragg mirror layers rather than within the cavity core.
@@
Figure 9: Transmittance and the resonant-mode field distribution of a Bragg cavity with the same parameters as in Fig. 5, loaded with an array of NCs ($R=150~\mathrm{nm}$), for periods of (a) $p=500~\mathrm{nm}$, (b) $p=750~\mathrm{nm}$, and (c) $p=900~\mathrm{nm}$. All structures are illuminated by a normally incident plane wave (along $y$-axis) with the $E$-field along $z$-axis (TM, $E_{z}$).
3.3 3.3 Tunability of the Bragg-Mie Modes Beyond the Size Parameter
@@
In addition to the conventional size parameters $x_{i}$ and $x_{e}$, which depend on wavelength, particle size and the refractive indices of the particle and surrounding medium,[Bohren2008] the coupled Bragg–Mie modes can be further tuned by external and structural parameters. Specifically, varying the angle of incidence (AoI) and the bare-cavity cutoff wavelength introduces additional means for controlling the resonance wavelength of the Bragg–Mie modes. These parameters enable tunability of the resonances beyond what is achievable with the size parameter alone, offering a richer degree of freedom for engineering their spectral position and field confinement. For instance, this approach can enable straightforward fine tuning and alignment of the mode resonances with the magnetic dipole resonances of doped atoms or quantum dots, facilitating enhancement or selective excitation of their intrinsic magnetic dipole transitions.
@@
Figure 10 illustrates this tunability for both the TE0-like (Fig. 10a) and TM1-like (Fig. 10b) modes by showing the transmittance of a doped Bragg cavity with the same parameters as in Fig. 5, for $p=500~\mathrm{nm}$ and different bare-cavity cutoff wavelengths of $\lambda_{c}=1450~\mathrm{nm}$, $1500~\mathrm{nm}$, and $1550~\mathrm{nm}$. For both modes, the resonance redshifts for larger cutoff wavelengths.
@@
Figure 10: Transmittance spectra of a doped Bragg cavity with period $p=500\,\mathrm{nm}$ for (a) TE0-like and (b) TM1-like modes at different cutoff wavelengths $\lambda_{c}=1450\,\mathrm{nm}$ (red dashed), $1500\,\mathrm{nm}$ (blue solid), and $1550\,\mathrm{nm}$ (green dotted). (c) TE-polarized transmittance of the same structure at $\lambda_{c}=1500\,\mathrm{nm}$ for incidence angles of $0^{\circ}$ (solid), $25^{\circ}$ (dashed), and $50^{\circ}$ (dash-dotted).
@@
Figure 10c shows the TE-polarized transmittance of the same structure at $\lambda_{c}=1500\,\mathrm{nm}$ for increasing AoI. The lower- and higher-energy modes shift and move closer together, illustrating an angle-tunable response; the lower-energy branch blue-shifts across the range, while the higher-energy branch shows a weaker, non-monotonic shift at larger angles.
@@
Figure 11 shows the TM-polarized transmittance of the same structure as in Fig. 10, but with $p=750~\mathrm{nm}$ and AoIs of $0^{\circ}$, $5^{\circ}$, and $15^{\circ}$. As seen in Fig. 10a, for non-zero AoIs, new resonances emerge with markedly high $Q$-factors, particularly at smaller angles (e.g., $Q\approx 1.5\times 10^{4}$ for AoI $5^{\circ}$, compared to $Q\approx 700$ for the TM1-like mode). This resonance is structurally similar to the TM1-like mode but exhibits a $90^{\circ}$ rotated field distribution. Upon increasing AoI, the mode redshifts while the TM1-like mode gradually vanishes. As illustrated in Fig. 11b for an AoI of $15^{\circ}$, additional resonances also appear at shorter wavelengths, with their fields predominantly distributed in the Bragg mirror layers rather than in the cavity core.
@@
Figure 11: TM-polarized transmittance of a Bragg cavity with $N=10$ pairs of cladding layers on each side of a half-wave core, designed for a cutoff wavelength of $1500~\mathrm{nm}$, and loaded with an array of cylindrical NCs ($R=150~\mathrm{nm}$) with period $p=750~\mathrm{nm}$. (a) Results for incidence angles of $0^{\circ}$ (dashed curve) and $5^{\circ}$ (solid curve). (b) Results for an incidence angle of $15^{\circ}$.
@@
The modes discussed above—including the TE- and TM-like resonances confined inside the NCs, the modes localized between the NCs, and the high-$Q$ resonance emerging at nonzero AoI in Fig. 11a—can all be identified as high-FoM NZI modes of the doped Bragg cavity. This is confirmed by the dispersion curves shown in Fig. 12a in the wavelength range of 1300–1650 nm, together with their corresponding transmission peaks presented in Fig. 12b.
@@
Figure 12: (a) Real and imaginary parts of effective refractive indices of TE and TM modes, together with their corresponding field distributions, in a Bragg cavity with $N=14$ pairs of cladding layers on each side of a half-wave core, loaded with an array of cylindrical NCs, with a period of $p=750~\mathrm{nm}$. (b) Transmission spectrum of the same cavity for TE (solid curve) and TM (dashed/dotted curves) modes.